# Joint Attention Mechanism Learning to Facilitate Opto-physiological Monitoring during Physical Activity

### v1

@@ Abstract
Opto-physiological monitoring is a non-contact technique for measuring cardiac signals, i.e., photoplethysmography (PPG). Quality PPG signals directly lead to reliable physiological readings. However, PPG signal acquisition procedures are often accompanied by spurious motion artefacts (MAs), especially during low-to-high-intensity physical activity. This study proposes a practical adversarial learning approach for opto-physiological monitoring by using a generative adversarial network with an attention mechanism (AM-GAN) to model motion noise and to allow MA removal. The AM-GAN learns an MA-resistant mapping from raw and noisy signals to clear PPG signals in an adversarial manner, guided by an attention mechanism to directly translate the motion reference of triaxial acceleration to the MAs appearing in the raw signal. The AM-GAN was experimented with three various protocols engaged with 39 subjects in various physical activities. The average absolute error for heart rate (HR) derived from the MA-free PPG signal via the AM-GAN, is 1.81 beats/min for the IEEE-SPC dataset and 3.86 beats/min for the PPGDalia dataset. The same procedure applied to an in-house LU dataset resulted in average absolute errors for HR and respiratory rate (RR) of less than 1.37 beats/min and 2.49 breaths/min, respectively. The study demonstrates the robustness and resilience of AM-GAN, particularly during low-to-high-intensity physical activities.
@@
[inst3]organization=School of Sport, Exercise and Health Sciences,addressline=Loughborough University, city=Loughborough, postcode=LE11 3TU, country=UK
[…]
[inst4]organization=Digital Environment Research (DERI),addressline=Queen Mary University of London, city=London, postcode=E1 1HH, country=UK
@@ 1 Introduction
Pulsatile waveform extracted from the PPG signal is a key component in obtaining physiological parameters, including heart rate (HR), respiration rate (RR) and oxygen saturation (SpO2). Obtaining reliable physiological parameters from PPG relies upon a high-quality signal. MAs, frequently generated during physical activity, impede the acquisition of clear PPG signals, thereby diminishing the stability and accuracy of vital sign measurements. MAs can arise due to alterations in the contact between the wearable sensor and the skin surface caused by movement, as well as from hemodynamic effects [1]. Various techniques have been investigated to extract MA to produce MA-reduced PPG signals. These methods include Empirical Mode Decomposition (EMD), Fast Fourier Transform (FFT), Independent Component Analysis (ICA), wavelet denoising [2], Adaptive Noise Cancellation (ANC) with Recursive Least Squares (RLS) and Adaptive-Size Least Mean Squares (AS-LMS) [3]. These approaches tend to underperform during medium-to-high-intensity physical activities, a challenge addressed by the TROIKA algorithm [4]. Also, such MA-removal methods are complex and a challenge for efficient embedding in a wearable system [5]. Although these studies adopt an acceleration reference for MA removal, the acceleration reference usually fails to precisely capture MA associated with subtle finger or wrist movements and gestures, leading to potentially uncorrelated MA behaviour. There has been a growing demand for additional, reliable physiological parameters beyond HR estimates, such as RR and SpO2, where obtaining a higher-quality pulsatile waveform is key. Consequently, additional PPG algorithms are required to extract relevant parameters/biomarkers during physical activities that are applicable for wearable opto-physiological monitoring. Such a prospect may well lie in the domain of Artificial Intelligence.
@@
Inspired by the rapid development of deep-learning (DL) techniques, several DL algorithms to extract physiological parameters from raw PPG signals have been recently presented. These techniques can be categorised into two types: end-to-end methods and feature extraction methods. The former directly establishes a mapping from raw PPG signals to the target physiological parameters [6], while the latter obtains the target parameters by extracting features pre-processed from raw PPG signals [7]. Data-driven, with strong fitting abilities, these neural networks often outperform traditional methods [8]. However, most DL end-to-end and feature extraction methods focus on extracting only the specific physiological parameter from input raw PPG signals, and often ignore the high-quality PPG signals extraction that dominates the reliable physiological parameter estimation.
@@
Some recent studies have demonstrated that using DL methods can benefit the noise and motion detection from the input noisy PPG signals [9]. Specifically, a one-dimensional (1D) Convolutional Neural Network (CNN) has been proposed to automatically learn the intrinsic features of the PPG signals and perform the required classification of ’clean’ or ’with artefacts’ [9]. Other DL methods transform PPG signals into two-dimensional (2D) images and employ a 2D CNN model for the signal quality classification on static and low-intensity exercise [10], [11]. Although these approaches have achieved high accuracy in HR evaluation, medium-to-high-intensity activities have not been evaluated, which is more challenging as MAs during these activities are more complex, with diverse morphologies. Attention mechanisms have emerged as a critical component in the advancement of sequence modelling and transduction models across a range of tasks, facilitating the modelling of dependencies irrespective of their positional distance within the input or output sequences [12]. Recent work combining generative adversarial network (GAN) and Attention has shown considerable promise: for instance, the study in [13] employed an attention-based generator, for the identification of local salient features, and incorporated dual discriminators to ensure the fidelity of the synthesised data across both time and frequency domains, in an attempt to generate synthetic ECG from PPG. A general issue with the method in [13], is that MAs are not to be considered. In contrast, a GAN was employed to eliminate low-intensity motion noise from PPG signals by transforming the PPG into a two-dimensional correlation image array, which enabled the use of a conventional image-based GAN architecture [11]. Similarly, the elimination of ocular artefacts from electroencephalography (EEG) data, crucial for numerous brain-computer interface applications, was explored in [14].
@@
To address these issues, this study presents an attention mechanism learning approach for robust PPG-based physiological parameters extraction by using an attention mechanism combined with the GAN for multi-sensor fusion to remove MAs from the raw PPG signals during low-to-high intensity physical activities. Pre-filtered, motion-contaminated PPG signals from an optoelectronic patch sensor, triaxial acceleration signals from an accelerometer sensor, and associated velocities were adopted as the inputs of the proposed AM-GAN model. Once the model learns how to map the contaminated PPG signals to MAs-free PPG signals, the generator generates the PPG signal similar to the ground-truth PPG signal, while the discriminator distinguishes the generated PPG signal from the ground-truth PPG signal. The attention mechanism is suitable for multi-sensor fusion to significantly enhance the generator’s ability to focus attention on the manner in which motion impacts the PPG waveform characteristics, thereby enhancing its capacity to eliminate MAs across a spectrum of physical activities ranging from low to high intensity. The contributions of this work are as follows:
@@
    To research how an Attention Mechanism learning suitable for multi-sensor fusion to be applied in a GAN network to effectively remove in-band and out-of-band noises.
@@
    To assess the ability of the attention-based GAN model to obtain a clean PPG pulsatile waveform by MA removal from raw PPG signal inputs using the outputs of the triaxial accelerometer in various scenarios of low-to-high intensity physical activities that the frequency range was shown to be between 0.1 and 4.0 Hz.
@@
    To validate the proposed AM-GAN model for extracting HR and RR for both intra-dataset and cross-dataset testing against other state-of-the-art (SOTA) methods.
@@
    To provide the benefit while utilising the attention mechanism to be justified with comparative output time and frequency-domain waveform, with and without the attention mechanism.
@@ 2 Methods and Materials
In this study, we propose an AM-GAN (depicted in Fig. 1) comprising an attention mechanism together with a GAN to learn the mapping from the MA-corrupted PPG signal to the MA-free PPG signal. Particularly, the proposed AM-GAN consists of three main parts: 1) Generator ($G$) together with the Attention Mechanism, 2) Discriminator ($D$), and (during training) 3) Motion removal algorithm.
During the training, the generator takes an MA-corrupted PPG signal $p(t)\in\mathcal{P}_{Tr}$ (where $\mathcal{P}_{Tr}$ denotes the training data set) as an input as well as triaxial acceleration signals $a_{x,y,z}(t)$ as a conditional input and generates an approximation to the MA-free PPG signal $s(t)=G(p(t),a_{x,y,z}(t))$ as its output. An attention mechanism is incorporated into the generator to force it to focus on features of the encoded PPG signal with the greatest importance. The discriminator is designed to distinguish the ground truth $s_{ref}(t)$ from the signals $s(t)=G(p(t),a_{x,y,z}(t))$. Finally, the motion removal algorithm, MR algorithm [15], extracts the reference PPG signals from MA-corrupted PPG signal $p(t)$, i.e., $s_{ref}(t)=MR(p(t))$, for training purposes.
To verify the performance of the trained network, the output of the generator is compared to the reference signal for signals in the testing data set, i.e., $p(t)\in\mathcal{P}_{Te}$. The physiological parameters (HR and RR) are also extracted from the MA-free PPG signals and compared with the ones from reference signals. The details of the proposed AM-GAN are described in the following subsections.
@@ 2.1 U-Net Generator Aligned with Attention Mechanism
The U-Net generator was initially proposed for image classification [16]. Fig. 2 shows the modified U-Net generator as a de-noising encoder-decoder with skip connections and the attention mechanism to take the noisy PPG signal as an input and the triaxial acceleration (ACC) as the conditional input. Meanwhile, the triaxial velocity (VEL) is created using the acceleration signal and constant matrix $Z$, i.e., $v_{x}=\frac{1}{f_{s}}Z[a_{x}]$, where $f_{s}$ is the sampling rate and $Z$ is the lower triangular matrix of constants (constant is 1).
In the modified U-Net generator architecture, 1D convolution and deconvolution functions are utilised instead of the typical 2D functions, and the latent vector $z$ is omitted, or more properly is replaced in essence by the motion reference proxy obtained from the triaxial ACC and VEL. These changes adapt the network better for processing 1D signals and help to minimise overfitting during the generation of the targeted PPG signals. Specifically, the encoding phases include four convolutional layers, each followed by batch normalisation (BN) and leaky rectified linear units (LeakyReLU). These components are integrated within a strided convolution process using a stride of two. The decoding process is the reversal of the encoding process, which is implemented through deconvolution layers followed by batch normalisation (BN) and rectified linear (ReLU) units to decode the de-noised PPG feature map culminating in the acquisition of the desired de-noised PPG signal. Additionally, the generator network employs skip connections that link each encoding layer to its corresponding decoding layer, effectively circumventing the central compression stage of the generator network. These skip connections enable the direct transfer of fine-grained features from the encoding stage to the decoding stage, addressing the potential loss of low-level details that can occur when all feature information pass through the central compression bottleneck of the network[16]. In the AM-GAN generator structure, the PPG signal encoding layer and motion reference encoding layer are concatenated to link the corresponding decoding layer.
@@ 2.2.1 Problem Formulation
Referring to the simplest version of our MR model [15], the noise $n[i]$, in the measured noisy PPG signal $p[i]$, can be approximated by a linear combination of acceleration outputs $a_{x}[i],a_{y}[i],a_{z}[i]$ and associated velocity (the integrated outputs $v_{x}[i],v_{y}[i]$ and $v_{z}[i]$) from the triaxial ACC, thus $n[i]=\boldsymbol{a}[i]\Xi$ where the columns of the $(N\times 6)$ array $\boldsymbol{a}[i]$are the six motion signals with coefficients given by the $(6\times 1)$ array $\boldsymbol{\Xi}$. An orthonormal basis for this subspace may be obtained from the six columns of the $(N\times 6)$ array $\boldsymbol{\Phi}=\boldsymbol{a}\boldsymbol{U}^{\intercal}\boldsymbol{D^{-1/2}}$, where $\boldsymbol{U}$ and $\boldsymbol{D}$ are the unitary matrix that diagonalises the $6\times 6$ array $\boldsymbol{a}^{\intercal}\boldsymbol{a}$, and the resulting diagonal form (i.e. $\boldsymbol{a}^{\intercal}\boldsymbol{a}=\boldsymbol{U}^{\intercal}\boldsymbol{D}\boldsymbol{U}$). The motion-suppressed PPG signal is then rewritten as the more obvious:
@@
where $\boldsymbol{\Phi}$ plays the role of a unit vector in standard Euclidean geometry. From (1), $\boldsymbol{s}^{\intercal}\boldsymbol{n}=0$, meaning that the method assumes that the clean signal and the noise are separable in the frequency domain. In the case that the spectra of $\boldsymbol{s}$ and $\boldsymbol{n}$ do overlap, (1) should remove some, but not necessary all, of the MAs; it will represent motion suppression rather than motion removal.

### v2

@@ Abstract
Opto-physiological monitoring including photoplethysmography (PPG) provides non-invasive cardiac and respiratory measurements, yet motion artefacts (MAs) during physical activity degrade its signal quality and downstream estimation concurrently. An attention-mechanism-based generative adversarial network (AM-GAN) was proposed to model motion artefacts and mitigate their impact on raw PPG signals. The AM-GAN learns how to transform motion-affected PPG into artefact-reduced waveforms to align with triaxial acceleration signals corresponding to artefact components gained from a triaxial accelerometer. The AM-GAN has been validated across four experimental protocols with 43 participants performing activities from low to high intensity (6–12km/h). With the public datasets, the AM-GAN achieves mean absolute error (MAE) for heart rate (HR) of 1.81 beats/min on IEEE-SPC and 3.86 beats/min on PPGDalia. On the in-house LU dataset, it shows the MAEs <1.37 beats/min for HR and 2.49 breaths/min for respiratory rate (RR). A further in-house C2 dataset with three oxygen levels (16%, 18%, and 21%) was applied in the AM-GAN to attain a MAE of 1.65% for SpO2. The outcome demonstrates that the AM-GAN offers a robust and reliable physiological estimation under various intensities of physical activity.
@@
[inst2]organization=School of Sport, Exercise and Health Sciences,addressline=Loughborough University, city=Loughborough, postcode=LE11 3TU, country=UK
@@ 1 Introduction
The pulsatile waveform extracted from the PPG signal is a key component in obtaining physiological parameters, including heart rate (HR), respiration rate (RR) and oxygen saturation (SpO2). Obtaining reliable physiological parameters from PPG relies upon a high-quality signal. Motion artefacts (MAs), frequently generated during physical activity, impede the acquisition of clear PPG signals, thereby diminishing the stability and accuracy of vital sign measurements. MAs can arise from changes in sensor, skin contact due to movement, as well as from haemodynamic effects [b3turcott2008hemodynamic]. Various techniques have been investigated to extract MA to produce MA-reduced PPG signals. These methods include Empirical Mode Decomposition (EMD), Fast Fourier Transform (FFT), Independent Component Analysis (ICA), wavelet denoising [b4pankaj2022review], Adaptive Noise Cancellation (ANC) with Recursive Least Squares (RLS), Adaptive-Size Least Mean Squares (AS-LMS) [b5ram2011novel] and an efficient envelope-based PPG denoising algorithm (EPDA) [b36bradley2024opening]. These approaches tend to underperform during medium-to-high-intensity physical activities, a challenge addressed by the TROIKA algorithm [b6zhang2014troika]. Also, such MA-removal methods are complex and a challenge for efficient embedding in a wearable system [b7nabavi2020robust]. Furthermore, a lightweight motion artefact intensity classification and removal (MAICR) framework is proposed to extract HR using acceleration time-domain information and PPG frequency-domain information [b34hao2024ppg]. Zhang et al. [b35zhang2024removal] proposed an MA-removal method combining CEEMDAN-based multiscale permutation entropy (CEEMDAN-MPE) with variable-step-size least mean squares (VS-LMS), effectively enhancing the accuracy of SpO2 measurement. Although these studies adopt an acceleration reference for MA removal, the acceleration reference usually fails to precisely capture MA associated with subtle finger or wrist movements and gestures, leading to potentially uncorrelated MA behaviour. There has been a growing demand for additional, reliable physiological parameters beyond HR estimates, such as RR and SpO2, where obtaining a higher-quality pulsatile waveform is key. Consequently, additional PPG algorithms are required to extract these additional parameters/biomarkers during physical activities that are advantageous during wearable opto-physiological monitoring. Such a prospect may well lie in the domain of Neural Networks.
@@
Inspired by the rapid development of deep-learning (DL) techniques, several DL algorithms to extract physiological parameters from raw PPG signals have been recently presented. These techniques can be categorised into two types: end-to-end methods and feature extraction methods. The former directly establishes a mapping from raw PPG signals to the target physiological parameters [b8luque2018end], while the latter obtains the target parameters by extracting features pre-processed from raw PPG signals [b9biswas2019cornet]. Data-driven, with strong fitting abilities, these neural networks often outperform traditional methods [b10esgalhado2021application]. However, most DL end-to-end based methods estimate specific physiological parameters directly from raw PPG, while neglecting the reconstruction of high-quality PPG waveforms that underpin reliable parameter estimation.
@@
Some recent studies have demonstrated that using DL methods can benefit the noise and motion detection from the input noisy PPG signals [b11goh2020robust]. Specifically, a one-dimensional (1D) convolutional neural network (CNN) has been used to learn intrinsic PPG features and classify signals as clean or artefact-contaminated [b11goh2020robust]. Other DL-based approaches convert PPG signals into two-dimensional (2D) representations and use 2D CNNs for signal-quality classification at rest and during low-intensity exercise. [b12chen2021signal], [b13afandizadeh2023accurate]. Although these approaches have demonstrated high accuracy in HR evaluation, their effectiveness during medium-to-high-intensity activities has not yet been assessed. Evaluating performance in such scenarios is crucial, as MAs during these activities are more complex and exhibit diverse morphologies. Currently, attention mechanisms have become central to modern sequence modelling and sequence-to-sequence tasks, enabling models to capture long-range dependencies regardless of position. [b30ashish2017attention]. Panagiotis et al. [b37kasnesis2023feature] proposed PULSE, a deep learning method leveraging temporal convolutions and multi-head attention to enhance HR monitoring. Additionally, Christodoulos et al. [b38kechris2024kid] introduced the adaptive Q-PPG [b31burrello2021q] framework, combining the baseline Q-PPG model with an attention mechanism to improve HR tracking accuracy. Recent work combining generative adversarial networks (GANs) with attention has shown promise. For example, [b31sarkar2021cardiogan] uses an attention-based generator to identify local salient features and employs dual discriminators to preserve fidelity in both the time and frequency domains when synthesising ECG from PPG. A general limitation of [b31sarkar2021cardiogan] is that MAs are not considered. In contrast, a GAN was employed to remove low-intensity motion artefacts from PPG signals by converting the signals into two-dimensional correlation images, enabling the use of a standard image-based GAN architecture [b13afandizadeh2023accurate]. Similarly, the elimination of ocular artefacts from electroencephalography (EEG) data, crucial for numerous brain-computer interface applications, was explored in [b17sawangjai2021eeganet]. Furthermore, Sawangjai et al. [b33sawangjai2025removal] proposed an attention GAN with dual discriminator to remove MAs without additional motion data from accelerometers or gyroscopes. However, their approach faces an obvious limitation in effectively addressing MAs during medium-to-high-intensity movements due to inherent architectural constraints.
@@
To tackle these challenges, the study proposes an attention mechanism-based learning approach capable of robustly extracting physiological parameters from multi-sensor data, specifically PPG and 3-axis acceleration signals, across multiple datasets. Hence, an attention mechanism combined with a GAN is chosen and utilised for multi-sensor fusion to effectively remove MAs from raw PPG signals during physical activities ranging from low to high intensity. Pre-filtered, motion-contaminated PPG signals from an optoelectronic patch sensor, triaxial acceleration signals from an accelerometer sensor, and associated velocities were adopted as the inputs of the proposed AM-GAN model. Once the model learns how to map the contaminated PPG signals to MA-free PPG signals, the generator generates a PPG signal similar to the ground-truth PPG signal. At the same time, the discriminator distinguishes the generated PPG signal from the ground-truth PPG signal. The attention mechanism is well-suited for multi-sensor fusion, significantly enhancing the generator’s capability to accurately identify how MAs the characteristics of PPG waveforms. Consequently, this improves the system’s effectiveness in removing MAs in physical activities ranging from low to high intensity. The relevant literatures are summarised in Appendix 1. The contributions of this work are outlined as follows:
@@
    We propose an updated attention mechanism integrated within a GAN specifically designed for multi-sensor fusion, enabling the effective removal of both in-band and out-of-band noise from PPG signals.
@@
    The capability of the attention-based GAN (AM-GAN) to generate clear and reliable PPG waveforms by removing MAs from raw signals using synchronised triaxial accelerometer data across various physical activities, ranging from low to high intensity within a frequency range of 0.1–4.0 Hz is evaluated.
@@
    The effectiveness of the proposed AM-GAN model is validated through the extraction of HR, RR, and SpO2. Comprehensive performance comparisons against state-of-the-art methods are conducted using both intra-dataset and cross-dataset evaluations.
@@
    The practical perspective is demonstrated while employing the attention mechanism by comparing outputs in both time and frequency domains, highlighting improvements achieved through the proposed approach.
@@ 2 Methods and Materials
In this study, we propose an AM-GAN (depicted in Fig. 1) comprising an attention mechanism together with a GAN to learn the mapping from the MA-corrupted PPG signal to the MA-free PPG signal. Particularly, the proposed AM-GAN consists of three main parts: 1) Generator ($G$) together with the Attention Mechanism, 2) Discriminator ($D$), and 3) Motion removal algorithm (during training).
During the training, the generator takes an MA-corrupted PPG signal $p(t)\in\mathcal{P}_{Tr}$ (where $\mathcal{P}_{Tr}$ denotes the training data set) as an input as well as triaxial acceleration signals $a_{x,y,z}(t)$ as a conditional input and generates an approximation to the MA-free PPG signal $s(t)=G(p(t),a_{x,y,z}(t))$ as its output. An attention mechanism guides the generator to focus on the most salient features in the encoded PPG signal. The discriminator is designed to distinguish the ground truth $s_{ref}(t)$ from the signals $s(t)=G(p(t),a_{x,y,z}(t))$. Finally, the motion removal algorithm, MR algorithm [b21zheng2023rapid], extracts the reference PPG signals from the MA-corrupted PPG signal $p(t)$, i.e., $s_{ref}(t)=MR(p(t))$, for training purposes.
To verify the performance of the trained network, the output of the generator is compared to the reference signal for signals in the testing data set, i.e., $p(t)\in\mathcal{P}_{Te}$. The physiological parameters (HR, RR, and SpO2) are also extracted from the MA-free PPG signals and compared with the ones from reference signals. The details of the proposed AM-GAN are described in the following subsections.
@@ 2.1 U-Net Generator Aligned with Attention Mechanism
The U-Net generator was initially proposed for image classification [b20ronneberger2015u]. Fig. 2 shows the modified U-Net generator as a de-noising encoder-decoder with skip connections and the attention mechanism to take the noisy PPG signal as an input and the triaxial acceleration (ACC) as the conditional input. Meanwhile, the triaxial velocity (VEL) is created using the acceleration signal and constant matrix $Z$, i.e., $v_{x}=\frac{1}{f_{s}}Z[a_{x}]$, where $f_{s}$ is the sampling rate and $Z$ is the lower triangular matrix of constants (constant is 1).
In the modified U-Net generator architecture, 1D convolution and deconvolution functions are utilised instead of the typical 2D functions, and the latent vector $z$ is omitted, or more properly is replaced in essence by the motion reference proxy obtained from the triaxial ACC and VEL. These changes adapt the network better for processing 1D signals and help to minimise overfitting during the generation of the targeted PPG signals. Specifically, the encoding phases include four convolutional layers, each followed by batch normalisation (BN) and leaky rectified linear units (LeakyReLU). These components are integrated within a strided convolution process using a stride of two. The decoding process is the reversal of the encoding process, which is implemented through deconvolution layers followed by batch normalisation (BN) and rectified linear (ReLU) units to decode the de-noised PPG feature map culminating in the acquisition of the desired de-noised PPG signal. Additionally, the generator network employs skip connections that link each encoding layer to its corresponding decoding layer, effectively circumventing the central compression stage of the generator network. These skip connections enable the direct transfer of fine-grained features from the encoding stage to the decoding stage, addressing the potential loss of low-level details that can occur when all feature information passes through the central compression bottleneck of the network[b20ronneberger2015u]. In the AM-GAN generator structure, the PPG signal encoding layer and motion reference encoding layer are concatenated to link the corresponding decoding layer.
@@ 2.2.1 Problem Formulation
Referring to the simplest version of our MR model [b21zheng2023rapid], the noise $n[i]$, in the measured noisy PPG signal $p[i]$, can be approximated by a linear combination of acceleration outputs $a_{x}[i],a_{y}[i],a_{z}[i]$ and associated velocity (the integrated outputs $v_{x}[i],v_{y}[i]$ and $v_{z}[i]$) from the triaxial ACC, thus $n[i]=\boldsymbol{a}[i]\Xi$ where the columns of the $(N\times 6)$ array $\boldsymbol{a}[i]$are the six motion signals with coefficients given by the $(6\times 1)$ array $\boldsymbol{\Xi}$. An orthonormal basis for this subspace may be obtained from the six columns of the $(N\times 6)$ array $\boldsymbol{\Phi}=\boldsymbol{a}\boldsymbol{U}^{\intercal}\boldsymbol{D^{-1/2}}$, where $\boldsymbol{U}$ and $\boldsymbol{D}$ are the unitary matrix that diagonalises the $6\times 6$ array $\boldsymbol{a}^{\intercal}\boldsymbol{a}$, and the resulting diagonal form (i.e. $\boldsymbol{a}^{\intercal}\boldsymbol{a}=\boldsymbol{U}^{\intercal}\boldsymbol{D}\boldsymbol{U}$). The motion-suppressed PPG signal is then rewritten as the more obvious:
@@
where $\boldsymbol{\Phi}$ plays the role of a unit vector in standard Euclidean geometry. From (1), $\boldsymbol{s}^{\intercal}\boldsymbol{n}=0$, meaning that the method assumes that the clean signal and the noise are separable in the frequency domain. In the case that the spectra of $\boldsymbol{s}$ and $\boldsymbol{n}$ do overlap, (1) should remove some, but not necessarily all, of the MAs; it will represent motion suppression rather than motion removal.